OpenStudy (anonymous):

I need help setting up a system of equations. A company sells nuts in bulk quantities. When bought in bulk, peanuts sell for $1.70 per pound, almonds for $2.60 per pound, and cashews for $3.25 per pound. Suppose a specialty shop wants a mixture of 240 pounds that will cost $2.75 per pound. Find the number of pounds of each type of nut if the sum of the number of pounds of almonds and cashews is three times the number of pounds of peanuts.

OpenStudy (jackellyn):

p=peanuts a=almonds c=cashews 1.70p + 2.60a + 3.25c = 2.75 a+c+p=240 a+c=3p ====> p= (a+c)/3

OpenStudy (anonymous):

thank you!